Using (x + a) (x + b) = x^2 + (a + b) x + ab, find
5.1 x 5.2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

5.1 x 5.2
= (5 + 0.1)(5 + 0.2)
= 5^2+ (0.1 + 0.2)5 + 0.1 x 0.2
= 25 + 1.5 + 0.02x
= 26.52

Thanks